Greg is twice as old as Judy. Three years ago, Greg was 3 times as old as Judy was then. Find their current ages.
Answer by JulietG(1812)

G = 2J [Greg is twice as old as Judy.]
G - 3 = 3(J -3) [Three years ago, Greg was 3 times as old as Judy was then. ]
.
Substitute the value of G in the first equation into the second equation.
(2J) - 3 = 3(J-3)
Distribute
2J - 3 = 3J - 9
Subtract 2J from each side
-3 = J - 9
Add 9 to each side
6 = J
.
If Judy is 6, and Greg is twice her age, then Greg is 12.
Let's plug it into the second equation to make sure it's correct.
G - 3 = 3(J -3)
12 - 3 = 3(6-3)
9 = 3(3)
9 = 9
Success!